# A Differentiable Model of Nucleic Acid Dynamics

### v1

@@ E.1.2 Worm-like-chain (WLC) Fitting
with $U^{tot}(X)=U^{ext}(X)+U(X)$ where $U(X)$ denotes the Hamiltonian of the DNA model parameterized by $\theta$ and $U^{ext}(X)$ denotes the energetic contribution corresponding to the external force. This yields a set of weights
@@ E.2 Stretch and Torsional Moduli
where $\langle\Delta\theta^{2}\rangle$ is the variance in deviation from equilibrium twist over some ensemble of samples and $L_{0}$ is the equilibrium end-to-end extension of the DNA. In the original presentation of oxDNA, $C$ was calculated using Monte Carlo moves chosen to enforce that the base pairs at the end of the central section remained perpendicular to the vector between their midpoints [18]. This allowed for an unambiguous definition of the axis about which torsion could be applied and twist measured. However, such MC moves are not implemented in a publicly available software package and such constraints cannot be applied as rigorously in MD, yielding poor estimates of C using Eqn. 59.
@@
Figure S7: Fitting to a target propeller twist distribution. A. An overview of the optimization problem. The initial distribution of propeller twists is Gaussian with $\mu=21.7^{\circ}$ and $\sigma^{2}=1.5$ and the objective is to minimize the KL-divergence with a target distribution. Here, we consider the case of increasing the variance ($\sigma^{2}=5.5$) while maintaining the mean ($\mu=21.7^{\circ}$). B. The progression of the mean, variance, and KL-divergence (i.e. the loss) throughout the optimization. C. Histograms of the sampled propeller twists with the initial parameters and the optimized parameters.
@@
where $\tilde{S}\equiv S-g^{2}/C$ is the effective extensional modulus. As in the case of the WLC fit and the passive calculation of $L_{ps}$, this entire calculation is made differentiable by reweighting the expectations computed in Equations 60 and 62 and using implicit derivatives for the line fitting procedures in Equations 63 and 64. In practice, a linear line fitting procedure has an analytical derivative.
@@
Figure S8: Fitting the sequence-specific stacking and hydrogen-bonding parameters to a target persistence length. A. The persistence length over time. B-C. The absolute change in the sequence-specific stacking and hydrogen bonding parameters between the initial and optimized values. We obey the same symmetries for sequence-specific parameters as described in Ref. [34]
@@ F.1 Duplex Melting Temperature Calculation
For a set of evenly-spaced temperatures $\{\tau_{lo},\cdots,\tau_{hi}\}$, we apply Equations 77-83 to compute $\{f_{\infty\text{, }\tau_{lo}},\cdots,f_{\infty\text{, }\tau_{hi}}\}$. Finally, we interpolate to find the temperature $\tau^{*}$ at which $f_{\infty\text{, }\tau^{*}}=0.5$. We can similarly interpolate to compute the width of the melting curve where

### v2

@@ E.1.2 Worm-like-chain (WLC) Fitting
with $U^{tot}(X)=U^{ext}(X)+U(X)$ where $U(X)$ denotes the Hamiltonian of the DNA model parameterized by $\theta$ and $U^{ext}(X)$ denotes the energetic contribution corresponding to the constant external force. This yields a set of weights
@@ E.2 Stretch and Torsional Moduli
where $\langle\Delta\theta^{2}\rangle$ is the variance in deviation from equilibrium twist over some ensemble of samples and $L_{0}$ is the equilibrium end-to-end extension of the DNA. In the original presentation of oxDNA, $C$ was calculated using Monte Carlo moves chosen to enforce that the base pairs at the end of the central section remained perpendicular to the vector between their midpoints [39]. This allowed for an unambiguous definition of the axis about which torsion could be applied and twist measured. However, such MC moves are not implemented in a publicly available software package and such constraints cannot be applied as rigorously in MD, yielding poor estimates of C using Eqn. 64.
@@
where $\tilde{S}\equiv S-g^{2}/C$ is the effective extensional modulus. As in the case of the WLC fit and the passive calculation of $L_{ps}$, this entire calculation is made differentiable by reweighting the expectations computed in Equations 65 and 67 and using implicit derivatives for the line fitting procedures in Equations 68 and 69. In practice, a linear line fitting procedure has an analytical derivative.
@@
Figure S8: Fitting the sequence-specific stacking and hydrogen-bonding parameters to a target persistence length. A. The persistence length over time. Iterations in which reference state are resampled are represented as scatter points. B-C. The absolute change in the sequence-specific stacking and hydrogen bonding parameters between the initial and optimized values. We obey the same symmetries for sequence-specific parameters as described in Ref. [23]
@@ F.1 Duplex Melting Temperature Calculation
For a set of evenly-spaced temperatures $\{\tau_{lo},\cdots,\tau_{hi}\}$, we apply Equations 82-88 to compute $\{f_{\infty\text{, }\tau_{lo}},\cdots,f_{\infty\text{, }\tau_{hi}}\}$. Finally, we interpolate to find the temperature $\tau^{*}$ at which $f_{\infty\text{, }\tau^{*}}=0.5$. We can similarly interpolate to compute the width of the melting curve where
@@ F.2 Hairpin Melting Temperature Calculation
Figure S10: Fitting oxDNA parameters to $S_{\text{eff}}$, $L_{ps}$, pitch, and $C$, with and without conflict-free updates. The target pitch is $10.25$ bp/turn, the target $S_{\text{eff}}$ is $1045$ pN, and the target $L_{ps}$ is $50$ nm, and the target $C$ is 436 $436\text{ pN}\cdot\text{nm}^{2}$. Outliers in the torsional modulus $C$ at iterations 53-55 with conflict-free updates are omitted and replaced with a dashed line. Iterations in which reference states are resampled are represented as scatter points.